OpenStudy (anonymous):

I KEEP GETTING y=(1/6)x+43/6 but it doesn't look right... is that the answer???

OpenStudy (jdoe0001):

that's what I got, yes

OpenStudy (jdoe0001):

hmnm wait... shoot misread my own writing lol

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

hmmm nope.... it's correct

OpenStudy (anonymous):

really?

OpenStudy (jdoe0001):

\(\bf slope =m = \cfrac{1}{6}\qquad\qquad (5, 8)\\ \quad \\ y-y_1=m(x-x_1)\implies y-8=\cfrac{1}{6}(x-5)\implies y-8=\cfrac{1}{6}x-\cfrac{5}{6}\\ \quad \\ y = \cfrac{1}{6}x-\cfrac{5}{6}+8\implies y = \cfrac{1}{6}x-\cfrac{5}{6}+\cfrac{48}{6}\implies y = \cfrac{1}{6}x+\cfrac{53}{6}\)

OpenStudy (jdoe0001):

woops, shoot, have a typo hold on

OpenStudy (jdoe0001):

\(\bf slope =m = \cfrac{1}{6}\qquad\qquad (5, 8)\\ \quad \\ y-y_1=m(x-x_1)\implies y-8=\cfrac{1}{6}(x-5)\implies y-8=\cfrac{1}{6}x-\cfrac{5}{6}\\ \quad \\ y = \cfrac{1}{6}x-\cfrac{5}{6}+8\implies y = \cfrac{1}{6}x-\cfrac{5}{6}+\cfrac{48}{6}\implies \color{blue}{y = \cfrac{1}{6}x+\cfrac{43}{6}}\)

OpenStudy (anonymous):

thanks :3

OpenStudy (jdoe0001):

yw